### v1

# A Cycle Ride to HDR: Semantics Aware Self-Supervised Framework for Unpaired LDR-to-HDR Image Translation

Abstract
[…]
1 Introduction
High Dynamic Range (HDR) [4, 57] images capture a broader range of intensity values compared to Low Dynamic Range (LDR) or Standard Dynamic Range (SDR) counterparts, which have a pixel bit depth of only $2^{8}$ intensity levels. From human vision perspective, LDR images are often not visually pleasing, while from computer vision applications perspective, they hold limited information.
Most commercial devices produce and display LDR content and some specialized hardware can support HDR, e.g., HDR cameras and sensors [37], which can capture images with more than $256$ intensity levels and HDR displays [12], which can display HDR intensity levels. Given that specialized hardware that captures HDR is expensive, researchers have been investigating approaches for accurate HDR reconstruction from LDR (i.e., inverse tone-mapping [78]) using conventional non-learning methods and modern data-driven approaches. For display purposes, however, one often needs to tone-map [27] the HDR content to LDR to fit the intensity range supported by standard hardware.
Early attempts on HDR reconstruction address tasks such as image enhancement, e.g., recovering missing information due to extreme lighting conditions [62] or low lighting conditions [84, 47] and compression [11]. Current applications of HDR reconstruction span multiple domains, including, robotics/machine vision [83], media and entertainment [29], gaming, mixed reality and novel view synthesis [72, 53, 75, 31], as well as medical imaging [31].
Early data-driven methods utilize Convolutional Neural Networks (CNN) [74], Transformers [73, 52] and Generative Adversarial Networks (GAN) [67, 63]. More recent approaches are based on Diffusion Models [15, 23], Neural Radiance Fields (NeRF) [31], and Gaussian Splatting [75]. Some methods use multi-exposed LDR [6, 71, 18, 48, 51, 45, 24, 90, 25, 69], while others use single-exposed LDR [5, 9, 88] paired with an HDR for supervised training.
Although these methods achieve excellent results in reconstructing HDR images, most of them require proper {LDR,HDR} paired datasets [7, 51, 19] for training. Consequently, the quality of the state-of-the-art data-driven HDR reconstruction methods depends on the quality of the available paired {LDR,HDR} datasets. There is a research gap in the field; on the one hand, the literature on unpaired HDR reconstruction is extremely limited [49], and on the other, only a few approaches utilize semantic and contextual information to guide the reconstruction process [23, 22, 52, 77].
To address this gap, we propose a method that leverages cycle consistency [89, 14] objective using unpaired data, where the model learns a mapping between domains, i.e., LDR $\leftrightarrow$ HDR. In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing CLIP [66] embeddings and loss based on semantic segmentation. The proposed method reconstructs artifact-free and visually impressive HDR from single-exposed LDR images. It also outperforms the most recent state-of-the-art which predominantly uses paired datasets for training. Our work makes the following contributions:
[…]
  • We propose novel generators based on modified U-Net architecture [70] that incorporates ConvLSTM-based feedback mechanism [30, 50] to mitigate visual artifacts in the HDR reconstruction (cf., Section 3.1).
  • We propose a CLIP embedding encoder for contrastive learning to minimize the semantic difference between LDR and reconstructed HDR pairs, while maximizing the difference between LDR and other {LDR,HDR} pairs (cf., Section 3.1).
  • We propose a novel loss function based on the Mean Intersection of Union (mIOU) metric to further ensure semantic consistency between the LDR and reconstructed HDR (cf., Section 3.2).
  • We perform thorough experimental validation for the contribution of all proposed components, both qualitatively and quantitatively (cf., Section 5).
2 Related Work
2.1 Non-Learning Approaches
The most popular approach in this category consists of multi-exposed LDR fusion to achieve high dynamic range in the output image [35]. This approach involves image feature alignment, calculating weights for feature mapping on the basis of image characteristics, and fusing the images on the basis of the weights to get the most appropriate exposures for each part of the image. Another approach is using histogram equalization [16] that allows for contrast and brightness levels re-adjustment in the over/underexposed regions of the image. Gradient domain manipulation [44] techniques enhance the granular details of an LDR image and expand its dynamic range. Pyramid-based image fusion [38] is a technique where image features are extracted into Laplacian pyramids [38] and then fused for the HDR. Some methods use Retinex-based [40, 58] tuning to effectively produce and display HDR images on consumer screens [40, 58]. Intensity Mapping Function based approaches [85, 17, 65] use transformations on the image pixels either by using Logarithmic mapping or Gamma correction to stretch the intensity range or enhance lower intensity pixels. The methods in this category, although time-efficient, face some issues including ghosting effects, artifacts, and blurring in the reconstruction, and they do not generalize well for variety of input LDR images.
2.2 Learning-Based Approaches
The methods in this category either use single-exposed LDR as input [71, 18, 48, 51, 45, 24, 90, 25] or alternatively, multi-exposed LDR [63, 9, 69, 88, 6]. Barua et al. [5] harnessed the concept of histogram equalization of input LDR images, in addition to the original LDR, to overcome the contrast and hue miss-representation in over/underexposed areas of the LDR image. Cao et al. [10] proposed a channel-decoupled kernel-based approach which combines the output HDR with the output of another architecture in a pixel-wise fashion. Luzardo et al. [54] proposed a method to enhance the artistic intent of the reconstructed HDR by enhancing the peak brightness in the reconstruction process. In addition, recently, Neural Radiance Fields (NeRF) [60, 31, 59, 53], Diffusion Models [15, 23], and Gaussian Splatting [75] have also been used to improve the reconstruction, both in terms of space and time complexity.
Some methods use weakly-supervised [45], self-supervised [88] and unsupervised [62] approaches for HDR reconstruction. Le et al. [45] proposed an indirect approach for HDR reconstruction from multi-exposed LDR using weak supervision. The method first outputs a stack of multi-exposed LDR, which are then merged using the state-of-the-art tool Photomatix [28] to obtain an HDR image. Zhang et al. [88] proposed a self-supervised approach for HDR reconstruction. The method requires multi-exposed LDR to learn the reconstruction network during training without the need for HDR counterparts. Nguyen et al. [62] proposed an unsupervised approach to recover image information from overexposed areas of LDR images. This approach does not require any HDR for supervision and instead uses pseudo ground truth images. Lee et al. [46] proposed a method with limited supervision for multi-exposed LDR generation consisting of a pair of networks that generate LDR with higher and lower exposure levels. Then the two exposure levels are combined to generate the HDR image.
GAN-Based Approaches. Generative Adversarial Networks (GAN) [21, 3] architectures are well-explored in this domain. Niu et al. [63] proposed a GAN-based method that can handle images with foreground motions by fusing multi-exposed LDR and extracting important information from the over/underexposed areas of the LDR. Raipurkar et al. [67] proposed a conditional GAN architecture that adds details to the saturated regions of the input LDR using a pre-trained segmentation model to extract exposure masks. Guo et al. [24] proposed a two-stage pipeline that extracts the over/underexposed features with high accuracy. GAN with attention mechanism first generates the missing information in those extreme exposure areas, and in the second stage, a CNN with multiple branches fuses the multi-exposed LDR from the previous stage to reconstruct the HDR. Nam et al. [61] proposed a GAN method that uses exposure values for conditional generation of multi-exposure stack that adapts well to varying color and brightness levels.
Li et al. [49] proposed a GAN-based method for unpaired multi-exposure LDR to HDR translation. The method introduces modified GAN loss and novel discriminator to tackle ghosting artifacts caused from the misalignment in the LDR stack and HDR. Wu et al. [82] proposed CycleGAN-like architecture for low-light image enhancement task using unpaired data for training. The method fuses the Retinex theory with CycleGAN concept to enhance the lighting conditions globally, recover color and reduce noise in the output HDR image.
Table 1: Summary of recent state-of-the-art methods. I/O: LDR used as input (SE: Single-exposed and ME: Multi-exposed), O/P: Reconstructs directly HDR (D) or multi-exposed LDR stack (I), Unpaired: Uses unpaired data, Context: Uses local/global image information and relationship among entities in the image, Semantics: Uses color/texture information and identity of the items in the image, Artifacts: Handles visual artifacts, TM: Also performs tone-mapping i.e. HDR $\rightarrow$ LDR.
Method I/O O/P Unpaired Context Semantics Artifacts TM
PSENet [62] SE D
SingleHDR(W) [45] SE I
UPHDR-GAN [49] ME D
SelfHDR [88] ME I
KUNet [77] SE D
Ghost-free HDR [52] ME D
DITMO [23] SE I
Ours SE D
Semantic and Knowledge-Based Approaches. Some methods attempt to use semantic/context information in the image or image formation process for HDR reconstruction. Wang et al. [77] proposed a method which approximates the inverse of the camera pipeline. Their knowledge-inspired block uses the image formation knowledge to address three tasks during HDR reconstruction: missing details recovery, image parameters adjustment, and image noise reduction. Goswami et al. [23] proposed a method to recover the clipped intensity values of an LDR/SDR image due to the tone-mapping process in the camera. The proposed method works in two stages - first it uses a semantic graph-based guidance to help the diffusion process with the in-painting of saturated image parts, and second, the problem is formulated as HDR in-painting from SDR in-painted regions. Liu et al. [52] proposed a vision transformer approach with context-awareness to remove ghosting effects in the output HDR. The model is a dual-branch architecture to capture both local and global context in the input image that enables the generation process to remove unwanted information in the output HDR and avoid artifacts.
Limitations. Our analysis reveals that methods based on single-exposed LDR fail to preserve image characteristics such as color hue and saturation, color contrast, and intensity brightness levels in the HDR. On the other hand, methods based on multi-exposed LDR produce unwanted visual artifacts, missing details and ghosting effects. While most methods produce excellent results in terms of visual quality, the main limitations of these approaches are the requirement of paired {LDR,HDR} datasets for training, and the lack of semantic and contextual knowledge guidance in the reconstruction process which could significantly improve the quality of output HDR images. Table 1 summarizes a comparison between recent state-of-the-art and our method in terms of various parameters.
Figure 1: Overview of the proposed method architecture (cf., Section 3.1) where $x$ and $y$ represent LDR and HDR images, respectively. The method is trained with five objectives: adversarial, cycle consistency, identity, contrastive, and semantic segmentation (cf., Section 3.2).
3 Method
[…]
3.1 Architecture Modules
We adopt a CycleGAN [89] architecture as the basis of our method depicted in Figure 1. The network includes two generators and two discriminators (cf., Supplementary for brief background on cycle consistency). Let us denote the two domains as $X$ for LDR images and $Y$ for HDR images. Furthermore, let us denote the generator used in the forward cycle that maps images from LDR to HDR as $G_{Y}$ and the one used in the backward cycle that maps images from HDR to LDR as $G_{X}$. Finally, let us denote the discriminator that discriminates between the reconstructed LDR and real LDR images as $D_{X}$ and the one that discriminates between the reconstructed HDR and real HDR images as $D_{Y}$. The images from the two domains can be represented as $\{x_{i}\}_{i=1}^{N}$ where $x_{i}\in X$ and $\{y_{j}\}_{j=1}^{M}$ where $y_{j}\in Y$. The data distribution of the two domains can be represented as $x\sim p_{\text{data}}(x)$ and $y\sim p_{\text{data}}(y)$.
Generators. The generators $G_{Y}$ and $G_{X}$ are based on U-Net architecture [70] that includes an encoder and a decoder block with skip connections from each level of the encoder to the decoder. In our generators, we propose a feedback mechanism [30, 86] between the encoder and decoder block. The rationale behind the feedback is to refine the features extracted from the encoder (during the first iteration of the feedback) to guide the decoder for better output image reconstruction over the rest of the iterations. Hence, the feedback block not only iterates over its own output but also re-runs the decoder situated ahead of it in each iteration while keeping the encoder frozen until the feedback iteration completes. The feedback is implemented with a ConvLSTM [30, 50] network and the number of iterations is fixed to $4$ based on an ablation experiment testing $2$, $3$, and $4$ iterations setups (cf., Supplementary for visual depiction of the proposed generators and implementation details).
Discriminators. The discriminators $D_{X}$ and $D_{Y}$ are based on [49, 32] (cf., Supplementary for implementation details).
Encoders. We also introduce a CLIP embeddings encoder $E$. Specifically we use the pre-trained Vision Transformer (ViT) [2] CLIP encoder from [66] which helps in extracting image embeddings with both local and global semantic context. For the forward cycle consistency we use $E(x)$ and $E(y)$. Then we add the embeddings from the encoder and feed them back to the bottleneck layer of the $G_{Y}$ decoder. Similarly, for the backward cycle consistency we add the embeddings from $E(y)$ and $E(x)$ and feed them back to the bottleneck layer of the $G_{X}$ decoder.
3.2 Loss Functions
We use five loss functions to train the proposed method. Two are standard CycleGAN loss functions, i.e., adversarial loss [20] and cycle consistency loss [89]. Inspired from [89] we also use an identity loss [76]. Finally, we introduce two novel loss functions for HDR reconstruction, i.e., contrastive and semantic segmentation loss.
The loss is calculated on tone-mapped versions of the reconstructed and real HDR images. The tone-mapping is performed on the basis of the $\mu$-law [34]. This is done to avoid the high intensity pixels of HDR images that can distort the loss calculation. The tone-mapping operator $T$ can be represented as follows:
$T(y_{j})=\frac{log(1+\mu y_{j})}{log(1+\mu)},$ (1)
where the amount of compression $\mu$ is 5000 following [39].
Adversarial Loss. We apply adversarial loss to both mappings. The mapping from LDR to HDR domain, i.e., $G_{Y}:X\rightarrow Y$ with the discriminator $D_{Y}$, can be expressed as:
$\begin{split}\mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)=\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\log D_{Y}(y)\right]+\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\log(1-D_{Y}(G_{Y}(x)))\right],\end{split}$ (2)
where $G_{Y}$ generates images that look similar to images from domain $Y$ while $D_{Y}$ distinguishes between generated samples $\hat{y}$ and real samples $y$. $G_{Y}$ aims to minimize this objective against $D_{Y}$ that aims to maximize it, i.e., $\min_{G_{Y}}\max_{D_{Y}}\mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)$.
The mapping from HDR to LDR domain, i.e., $G_{X}:Y\rightarrow X$ with the discriminator $D_{X}$, can be expressed as:
$\begin{split}\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)=\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\log D_{X}(x)\right]+\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\log(1-D_{X}(G_{X}(y)))\right],\end{split}$ (3)
where $G_{X}$ generates images that look similar to images from domain $X$ while $D_{X}$ distinguishes between generated samples $\hat{x}$ and real samples $x$. Similar to above, $G_{X}$ aims to minimize this objective against $D_{X}$ that aims to maximize it, i.e., $\min_{G_{X}}\max_{D_{X}}\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)$.
Cycle Consistency Loss. The adversarial loss does not guarantee learning without contradiction i.e., the forward $G_{Y}:X\rightarrow Y$ and backward $G_{X}:Y\rightarrow X$ mappings might not be consistent with each another. Hence, we also incorporate a cycle consistency loss to prevent mutual contradiction of the learned mappings $G_{Y}$ and $G_{X}$. For each image $x_{i}$ from the LDR domain, the cycle of reconstruction i.e., from LDR to HDR and then back to LDR must result back in the original image $x_{i}$. Hence, we can define the forward cycle consistency as: $x_{i}\rightarrow G_{Y}(x_{i})\rightarrow G_{X}(G_{Y}(x_{i}))\approx x_{i}$. Similarly, the backward cycle consistency can be represented as: $y_{j}\rightarrow G_{X}(y_{j})\rightarrow G_{Y}(G_{X}(y_{j}))\approx y_{j}$. The loss can be formulated as:
$\begin{split}\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})=\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(G_{Y}(x))-x\|_{1}\right]+\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(G_{X}(y))-y\|_{1}\right],\end{split}$ (4)
where $\|\cdot\|_{1}$ is the L1 norm (cf., Supplementary for visual depiction of the cycle consistency loss).
Identity Loss. For LDR $\leftrightarrow$ HDR translation tasks, we also find that adversarial and cycle consistency formulations alone cannot preserve the color and hue information. This is due to incorrect mapping of color shades from LDR to HDR domains by the generators stemming from the underlying difference in dynamic ranges. Therefore, we force the generators to replicate an identity mapping by providing target domain images. This loss can be expressed as:
$\begin{split}\mathcal{L}_{\text{id}}(G_{Y},G_{X})=\\ \mathbb{E}_{y\sim p_{\text{data}}(y)}\left[\|G_{Y}(y)-y\|_{1}\right]+\\ \mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\|G_{X}(x)-x\|_{1}\right].\end{split}$ (5)
Contrastive Loss. This loss is based on embeddings extracted using a CLIP encoder and ensures semantic information preservation across domains. Here, we do not directly extract the embeddings from LDR images but instead, use a histogram-equalized version processed using the OpenCV function equalizeHist [8]. Histogram equalization improves the pixel visibility in extreme lighting areas or areas with shadows/darkness of an image by re-adjusting the contrast and saturation levels. This is done by spreading out the frequent pixel intensity values across $256$ bins. Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image. For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(y)$ we first define the cosine similarity between them as follows:
$\text{sim}(\bar{x},\bar{y})=\frac{\bar{x}\cdot\bar{y}}{\|\bar{x}\|\|\bar{y}\|},$ (6)
[…]
$\begin{split}\mathcal{L}_{\text{con}}=-\frac{1}{N}\times\\ \sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\bar{x}_{i},\bar{y}_{i})/\tau)}{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(\exp(\text{sim}(\bar{x}_{i},\bar{x}_{j})/\tau)+\exp(\text{sim}(\bar{x}_{i},\bar{y}_{j})/\tau))},\end{split}$ (7)
where $N$ is the batch size, $\exp$ represents the exponential function, and $\tau$ represents the temperature parameter which controls the amount of emphasis given in distinguishing between positive and negative pairs. This loss is calculated in both the forward and backward cycles with the input and output being swapped. This loss replicates the contrastive learning paradigm of CLIP for {image,image} instead of {image,text} pairs (cf., Supplementary for visual depiction of the contrastive loss and an ablation experiment for $\tau$).
Semantic Segmentation Loss. This loss is based on segmentation masks. The Mean Intersection of Union (mIoU) metric measures the amount of overlap between ground truth and predicted segmentation masks. Similar to the previous loss function, we use histogram-equalized versions of the LDR images processed using the OpenCV function equalizeHist [8]. We choose equalized images instead of original LDR because segmentation on low-light or extreme brightness images does not yield good results. We use Segment Anything (SAM) [42] to generate segmentation classes in the histogram-equalized LDR and reconstructed tone-mapped HDR images (cf., Supplementary for visual depiction of the semantic segmentation loss). This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images. We can define the IoU metric as:
$\text{IoU}_{c}=\frac{x_{c}\cap y_{c}}{x_{c}\cup y_{c}},$ (8)
where $x_{c}$ and $y_{c}$ represent the segmentations for class $c$ in image $x$ and $y$, respectively. $\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$. The mean IoU over all segmentation classes can be formulated as:
$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (9)
We represent the semantic segmentation loss as:
$\mathcal{L}_{\text{sem}}=1-\text{mIoU}.$ (10)
$\mathcal{L}_{\text{full}}=\mathcal{L}_{\text{GAN}}(G_{Y},D_{Y},X,Y)+\mathcal{L}_{\text{GAN}}(G_{X},D_{X},Y,X)+\lambda(\mathcal{L}_{\text{cyc}}(G_{Y},G_{X})+0.5\times\mathcal{L}_{\text{id}}(G_{Y},G_{X}))+\alpha\mathcal{L}_{\text{con}}+\beta\mathcal{L}_{\text{sem}}$ (11)
Final Loss. The full objective is expressed in Eq. (11), where $\lambda$ scales the relative importance of the cycle consistency and identity loss. $\lambda$ is set to $10$ in our experiments and the weight for identity loss is $0.5$ inspired from the setup in the original cycle consistency work [89]. $\alpha$ and $\beta$ are weights for the contrastive and semantic segmentation loss, respectively. Both values are set to $2$ (cf., Supplementary for ablation experiments for $\alpha$ and $\beta$).
4 Experiments
4.1 Implementation
The proposed method is implemented in PyTorch [^1] https://pytorch.org/ on Ubuntu 20.04.6 LTS workstation with Nvidia Quadro P5000 GPU with 16GB memory, Intel® Xeon® W-2145 CPU at 3.70GHz with 16 CPU cores, 64GB RAM, and 2.5TB SSD. We used Adam optimizer [41] to train the models for $170$ epochs. Batch size of $1$ is expected to work well in cycle consistency models [89] however, since we are using contrastive loss on positive and negative pairs, the batch size is set to $4$. We used learning rate of $4\times 10^{-4}$ for the generators and $2\times 10^{-4}$ for the discriminators. The learning rate is kept constant for the first $100$ epochs and subsequently set to decay linearly to 00. For all models, we resized the input images to $512\times 512$. The HDR images used in the text are tone-mapped with Reinhard’s operator [68].
4.2 Datasets
For the primary comparison of our method with the state-of-the-art, we consider the HDRTV [13], NTIRE [64], and HDR-Synth & HDR-Real [51] paired datasets. HDRTV has $1235$ samples for training and $117$ for testing. The NTIRE dataset consists of approximately $1500$ training, $60$ validation, and $201$ testing samples. HDR-Synth & HDR-Real dataset consists of $20537$ samples. Other paired datasets used in our evaluations are DrTMO [19] ($1043$ samples), Kalantari [36] ($89$ samples), HDR-Eye [43] (46 samples), and LDR-HDR Pair [33] ($176$ samples). We choose these datasets because they consist of a balance between real and synthetic images as well as image and scene diversity. We perform $80/20$ split for training and testing unless specified otherwise. For methods working with single-exposed LDR inputs, we use only one LDR from datasets with multi-exposed LDR. For methods working with multi-exposed LDR inputs, we generate the required exposures using the OpenCV function convertScaleAbs [8] for datasets with only single-exposed LDR images.
4.3 Metrics
We use four metrics to report the results. High Dynamic Range Visual Differences Predictor (HDR-VDP-2) [55] or Mean Opinion Score Index (Q-Score) is used for evaluation replicating the human vision model. Structural Similarity Index Measure (SSIM) [79, 80, 81] is used to compare block-wise correlations on the basis of structural similarity, luminance, and contrast information. Peak Signal-to-Noise Ratio (PSNR) [26] (in dB) is used for a pixel-to-pixel comparison for noise in the signals [^2] PSNR is computed on $\mu$-law based tone-mapped HDR images.. Learned Perceptual Image Patch Similarity (LPIPS) [87] is used to compare between the high-level features in the images, such as the semantic entities present in the scene, and evaluates our semantic and contextual knowledge based contributions. This metric aligns with the perceptual judgement of humans.
Table 2: HDR reconstruction (inverse tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with supervised (gray) and unsupervised/weakly-supervised/self-supervised (black) learning methods trained on the paired datasets HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51].
[…]
HDRCNN [18] 20.12 0.724 0.422 60.12
DrTMO [19] 20.43 0.711 0.412 60.81
ExpandNet [56] 21.44 0.854 0.451 58.12
FHDR [39] 20.84 0.841 0.307 62.45
SingleHDR [51] 21.01 0.851 0.402 64.21
Two-stage [1] 34.29 0.856 0.287 61.78
HDR-GAN [63] 30.11 0.912 0.411 65.01
KUNet [77] 37.21 0.974 0.051 62.14
Ghost-free [52] 40.32 0.966 0.043 63.02
ArtHDR-Net [6] 37.12 0.961 0.321 63.43
HistoHDR-Net [5] 38.21 0.968 0.301 65.15
UPHDR-GAN [49] 39.98 0.965 0.044 63.54
PSENet [62] 27.35 0.856 0.274 62.89
SingleHDR(W) [45] 30.79 0.952 0.055 62.95
SelfHDR [88] 39.51 0.972 0.037 64.77
Ours 40.01 0.979 0.021 68.41
Table 3: HDR reconstruction (inverse tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with UPHDR-GAN trained on unpaired datasets.
[…]
UPHDR-GAN [49] 42.51 0.980 0.043 66.13
Ours 43.92 0.989 0.020 69.32
5 Results
5.1 HDR Reconstruction
For an extensive comparison we choose a combination of methods that utilize different learning paradigms including, HDRCNN [18], DrTMO [19], ExpandNet [56], FHDR [39], SingleHDR [51], Two-stage [1], KUNet [77], HistoHDR-Net [5], HDR-GAN [63], ArtHDR-Net [6], Ghost-free HDR [52], SelfHDR [88], UPHDR-GAN [49], SingleHDR(W) [45], and PSENet [62]. These state-of-the-art approaches include methods with single-exposed and multi-exposed LDR inputs, direct and indirect (i.e., either directly reconstructing HDR images or generating multi-exposed LDR stacks), GAN-based, Diffusion-based, Transformer-based, unsupervised/weakly-supervised/self-supervised, unpaired learning, and knowledge-conditioned.
Table 4: LDR reconstruction (tone-mapping) learned with our self-supervised learning approach. Quantitative comparison with the state-of-the-art one-mapping operators.
[…]
$\mu$-law operator [34] 28.68 0.891 0.311
Reinhard’s operator [68] 30.12 0.903 0.342
Photomatix [28] 32.11 0.922 0.136
Ours 31.71 0.941 0.092
Table 2 summarizes the results of evaluation on HDRTV [13], NTIRE [64], HDR-Synth & HDR-Real [51] paired datasets, the most widely used datasets by the state-of-the-art (cf., Supplementary for an extensive qualitative evaluation). In the top part of the table, we compare our method with $11$ supervised learning approaches. The proposed method outperforms those approaches in terms of SSIM, LPIPS, and HDR-VDP-2. In terms of PSNR, it is second best after Ghost-free HDR. In the bottom part of the table, we compare our method with $4$ unsupervised/weakly-supervised/self-supervised learning methods. The proposed method outperforms those approaches on all evaluation metrics (cf., Supplementary for visual depiction of the quantitative HDR reconstruction results). It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training. However, in all experiments, our method is trained in a self-supervised fashion using only unpaired samples (cf., Section 3).
To further demonstrate the strength of our approach we extended the train set used in the previous experiment with: 1) The HDR images from DrTMO [19], Kalantari [36], HDR-Eye [43], LDR-HDR Pair [33], and GTA-HDR [7] ($1500$ random samples), i.e., additional $2854$ HDR images; and 2) LDR images from GTA-HDR [7] ($2000$ random samples) and $50\%$ of the LDR images in DrTMO, Kalantari, HDR-Eye, and LDR-HDR Pair, i.e., additional $2677$ LDR images. We kept the test set identical to the previous experiment in order to compare with the state-of-the-art paired data based methods. Given that the only other method that supports this type of unpaired train data is UPHDR-GAN, we perform a direct comparison with that approach in Table 3. The results demonstrate that our method outperforms UPHDR-GAN and sets new state-of-the-art on all metrics (cf., Supplementary for more unpaired data experiments).
Table 5: Quantitative evaluation of our feedback U-Net-based generators for HDR reconstruction. For evaluation we use the dataset employed by SingleHDR(W), i.e., DrTMO [19]. The models marked with “(mod)” use feedback U-Net-based generators and the rest use the same U-Net generator as SingleHDR(W).
[…]
SingleHDR(W) [45] 30.79 0.952 0.055 62.95
SingleHDR(W) (mod) 30.03 0.961 0.036 63.91
Ours 35.11 0.969 0.029 67.88
Ours (mod) 43.01 0.988 0.023 68.21
Table 6: Architecture ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets. Uo: The U-Net generator used in SingleHDR(W) [45], Uf: The proposed U-Net generator, E: CLIP embedding encoder.
[…]
CycleGAN [89] 21.34 0.849 0.351 61.55
Uo 25.13 0.851 0.341 61.79
Uf 24.07 0.879 0.126 63.16
Uf+E 29.19 0.885 0.122 64.21
Uf+E+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
Uf+E+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
Uo+E+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 35.15 0.953 0.042 67.18
Ours 40.01 0.979 0.021 68.41
5.2 LDR Reconstruction
For a complete evaluation, we also report results for LDR image reconstruction in Table 4 and compare with the state-of-the-art tone-mapping operators $\mu$-law [34], Reinhard’s [68], and Photomatix [28]. The test data is identical to the one used in the HDR reconstruction experiments. The results demonstrate that our method outperforms the other approaches in terms of SSIM and LPIPS. In terms of PSNR, it is second best after Photomatix [68] (cf., Supplementary for more LDR reconstruction experiments).
5.3 Ablation Results
Generators. We propose a novel generator that modifies the U-Net architecture by introducing a feedback mechanism (cf., Section 3.1). The feedback helps in avoiding artifacts in image-to-image translation tasks where huge amounts of training data might make the process of hallucinating details difficult. To test this hypothesis, we first replaced the original U-Net used in the SingleHDR(W) [45] method with our U-Net generator. We see that leads to improvement in all metrics as shown in Table 5. The table also demonstrates the improvement in our model when we use the feedback U-Net instead of the original U-Net of SingleHDR(W) (cf., Supplementary for qualitative results in support of this quantitative evaluation).
Architecture. Our method uses the cycle consistency concept of CycleGAN and introduces many new components to address the LDR $\leftrightarrow$ HDR translation task. The most significant components are the CLIP embedding encoder and feedback U-Net generators (cf., Section 3.1). We perform a thorough ablation on each of the components to highlight the contributions. Table 6 summarizes the results where we add each component to the architecture as we go down. The first row lists the results from the original CycleGAN method. The second row is the CycleGAN utilizing the U-Net generator from SingleHDR(W) [45]. The third row is the CycleGAN utilizing our feedback U-Net generator. This results in a significant improvement in SSIM, LPIPS, and HDR-VDP-2. The fourth row lists the results when we add the CLIP embedding encoder to the architecture. This results in an improvement in all metrics. The fifth and sixth rows depict the improvements after adding contrastive and semantic segmentation objectives in the architecture (cf., Section 3.2). The seventh row depicts the results with all components in our architecture but with the original U-Net from SingleHDR(W). Finally, the last row is the full architecture with all components which achieves the best performance on all metrics.
Loss Functions. We also study the loss functions and their contribution towards the overall results keeping the architectural components constant. Table 7 summarizes the results. The first row is the results from the standard adversarial and cycle consistency loss. In the second row we add the identity loss which improves all metrics (cf., Supplementary for qualitative results in support of $\mathcal{L}_{\text{id}}$). In the third and forth rows we add the contrastive and semantic segmentation loss. This results in an improvement in all metrics (cf., Supplementary for qualitative results in support of $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{sem}}$). Finally, adding all loss functions achieves the best results on all the metrics.
Table 7: Loss functions ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
[…]
$\mathcal{L}_{\text{GAN}}$+$\mathcal{L}_{\text{cyc}}$ 24.55 0.878 0.125 62.11
+$\mathcal{L}_{\text{id}}$ 29.19 0.885 0.122 64.21
+$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{con}}$ 36.45 0.911 0.047 66.56
+$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{sem}}$ 34.21 0.897 0.094 66.12
+$\mathcal{L}_{\text{id}}$+$\mathcal{L}_{\text{con}}$+$\mathcal{L}_{\text{sem}}$ 40.01 0.979 0.021 68.41
We also perform separate tests on contrastive and semantic segmentation loss to examine the contribution of histogram-equalized LDR compared to original LDR images. We summarize these results in Tables 8 and 9. In both cases the loss calculated between the histogram-equalized versions of the LDR (instead of original the LDR) and the recnstructed HDR is better. This is due to the fact that histogram-equalized LDR can reveal semantic information in extremely over/underexposed areas of the original LDR images which in turn improves the extracted CLIP embeddings for $\mathcal{L}_{\text{con}}$ and segmentation maps for $\mathcal{L}_{\text{sem}}$.
Table 8: Contrastive loss ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
[…]
LDR 39.12 0.966 0.030 67.11
LDRHist 40.01 0.979 0.021 68.41
Table 9: Segmentation loss ablation results with the HDRTV [13], NTIRE [64] and HDR-Synth & HDR-Real [51] datasets.
[…]
LDR 38.89 0.956 0.033 66.92
LDRHist 40.01 0.979 0.021 68.41
6 Conclusions
Most current state-of-the-art methods for HDR reconstruction (inverse tone-mapping) require high-quality paired {LDR,HDR} datasets for model training. Consequently, the quality of the state-of-the-art in inverse tone-mapping depends on the quality of the available paired {LDR,HDR} datasets. Additionally, only a few approaches utilize semantic and contextual knowledge guidance in the reconstruction process which could significantly improve the quality of output HDR images.
[…]